OpenStudy (anonymous):

Stuck =| Please Help ! Thanxx =) Solve x – 2y = 6 3x – 6y = 18

OpenStudy (anonymous):

they are the same line, so you cannot find one point of intersection, they intersect everywhere

OpenStudy (anonymous):

they overlap if you graph them and again x and y can be all real numbers

OpenStudy (anonymous):

if you multiply the first equation all the way across by 3 you will see you get the second equation

OpenStudy (anonymous):

@Satellite73: So It WOuld Be DOne Like This : 3(x - 2y = 6) 3x - 6y = 18 Riqht ??

OpenStudy (anonymous):

3x - 6y = 18 -3x -3x = -6y = -3x + 18

OpenStudy (anonymous):

yes you get the same line

OpenStudy (anonymous):

no point in trying to solve it once you see they are the same

OpenStudy (anonymous):

once you write \[3(x - 2y = 6)\] \[ 3x - 6y = 18\] you see that the first line and the second line are really the same line